# Fourier dimension at $C^{1+\alpha}$ regularity

### v2

@@ Abstract.
We prove power Fourier decay for equilibrium states of nonlinear $C^{1+\alpha}$ iterated function systems, giving first examples of truly $C^{1+\alpha}$ IFSs with attractors of positive Fourier dimension that are not conjugated to a linear IFS previously out of reach of existing techniques. The key new ingredient is a Measure Non-Linearity (MNL) criterion tailored to $C^{1+\alpha}$ IFSs, coupling oscillations in geometric and symbolic directions and allowing us to prove the necessary multiscale non-linearity condition for Fourier decay via an induction-on-scales scheme. This bypasses the classical obstructions that require either large dimension enabling stationary phase comparisons (Kaufman measures), or smoothness or cohomological conditions on the roof function $\tau$ such as $\tau\in C^{1}$ (Dolgopyat’s method), $\tau\in C^{\alpha}$ but with smooth underlying dynamics (Tsujii–Zhang strategy), or non-flat $C^{\alpha}$ conjugacy to a constant (fractal van der Corput lemma). The method also applies to parabolic systems including limit sets of geometrically finite groups, Manneville-Pommeau maps, and Lorenz-type maps.
@@ 1. Introduction
1.1. Background and summary of the main results
The study of Fourier decay of dynamically defined measures has seen rapid progress in recent years. Given a probability measure $\mu$ on $\mathbb{R}$, we are interested in the high-frequency asymptotics of its Fourier transform
$\widehat{\mu}(\xi)=\int e^{-2\pi i\xi x}\,d\mu(x),\qquad\xi\in\mathbb{R},$
with particular emphasis on singular measures $\mu$ arising in hyperbolic and parabolic dynamics, fractal geometry, and geometric group theory. Examples include self-conformal measures, equilibrium states for uniformly and non-uniformly hyperbolic maps and flows, and Patterson–Sullivan measures for isometric group actions generating fractal limit sets. Fourier decay for such stationary measures for chaotic dynamical systems plays a central role in several areas, including bounding resonance free regions for scattering resonances in quantum chaos [13, 22], exponential mixing of flows and the decay of correlations (e.g. [31, 45]), equidistribution and Diophantine properties of numbers when applied to Gauss map [39], dispersive estimates for quasiperiodic Schrödinger operators [26] and stable intersections and homoclinic tangencies [7]. We refer to [40] for a broader survey on the recent developments.
Most of the literature on Fourier decay in dynamical settings has focused on systems with at least $C^{2}$ regularity or analyticity, where powerful tools such as Dolgopyat’s method [21] apply. This method, and its later refinements, exploits smooth holonomies and the structure of hyperbolic codings with finite alphabet to establish exponential mixing and Fourier decay [25, 13, 35, 2, 4, 41, 8, 10, 27, 28]. However, many natural dynamical systems of interest have only $C^{1+\alpha}$ regularity, such as codings of hyperbolic diffeomorphisms on basic sets due to $C^{1+\alpha}$ holonomies [44, 26], or exhibit parabolic features leading to infinite alphabet codings, including Manneville–Pommeau maps [18], Bunimovich billiards [16, 34, 14], geodesic flows on surfaces of non-positive curvature and Lorenz flows [33, 19], or hyperbolic surfaces with cusps [31].
The recent breakthrough of Tsujii and Zhang [44] on exponential mixing of 3D Anosov flows extended Dolgopyat’s method to rougher holonomies via new UNI conditions, and was adapted in [26] to prove power Fourier decay for equilibrium states of various Axiom $A$ diffeomorphisms, e.g. ones coming from the trace map of the Fibonacci Hamiltonian. However, these methods still fundamentally depend on the underlying smoothness of the dynamics (as there the roof function is smooth and one compares the temporal distance function to polynomials) and do not apply to equilibrium states of general $C^{1+\alpha}$ IFSs where both the roof and temporal distance are not smooth.” See Section 1.2 for further discussion comparing to [44, 26].
In this paper we introduce an induction-on-scales approach at the $C^{1+\alpha}$ regularity threshold that allows us to prove power Fourier decay for truly $C^{1+\alpha}$ IFSs. The method relies on verifying a multiscale nonlinearity condition Quantitative Non-Linearity (QNL) which yields power Fourier decay for oscillatory integrals with slowly varying amplitudes and phases (Theorem 1.1). In $C^{2}$ regularity, this is already more or less well-established route originating from the work of Bourgain and Dyatlov [13] and series of works after that [30, 41, 27, 10, 9, 26]. However, the main novelty of our work are the ways we verify QNL in rough settings where existing techniques do not apply. We develop two complementary routes which both base on induction-on-scales.
First, in the genuinely $C^{1+\alpha}$, non-conjugated IFS setting we introduce Measure Non-Linearity (MNL), a condition that couples symbolic and geometric oscillations; using an induction-on-scales tree scheme that avoids spectral gaps for twisted transfer operators, we show MNL implies QNL in the finite-branch case (Theorem 1.3). Then via a fixed-point argument on a moduli space of IFSs that may be of independent interest, we construct explicit truly $C^{1+\alpha}$ examples with power Fourier decay (Theorem 1.4). We expect this approach to apply more broadly in low-regularity dynamical systems. Second, in smoother categories we prove that uniform non-integrability (UNI) implies QNL (Theorem 1.2) via a $C^{\alpha}$ tree argument rather than Dolgopyat estimates. Because the framework is compatible with Young towers, this yields parabolic applications such as Manneville–Pommeau and Lorenz models, and via Li–Pan’s coding [31] gives Fourier decay for Patterson–Sullivan measures on geometrically finite hyperbolic surfaces with cusps (Theorems 1.5 and 1.6). Let us now describe our results in detail in the following Sections 1.2 and 1.3.2.
1.2. Results in $C^{1+\alpha}$ regularity
To state our results, we will need some notation from symbolic dynamics and thermodynamical formalism of subshifts of iterated function systems. The framework is that of equilibrium states for (possibly infinitely branched) uniformly hyperbolic $C^{1+\alpha}$ Markov systems with bounded distortions. This includes finite-branch $C^{1+\alpha}$ hyperbolic systems (by allowing equilibrium states that assign zero mass to the tail), and it is chosen specifically so as to also cover parabolic examples via Young tower constructions and codings that represent a finite-branch parabolic system as an infinite-branch hyperbolic IFS.
Symbolic coding. Let $\{I_{a}\}_{a\in\mathcal{A}}$ be pairwise disjoint open intervals in $[0,1]$, where $\mathcal{A}$ is finite or countable, and set $I=\bigcup_{a\in\mathcal{A}}I_{a}$. Suppose $T:\overline{I}\to\mathbb{R}$ is a $C^{1}$ map with the Markov property: whenever $T(I_{b})\cap I_{a}\neq\emptyset$, we have $T(\overline{I_{b}})\supset\overline{I_{a}}$. Define the transition matrix $A$ by $A(a,b)=1$ if $T(\overline{I_{a}})\supset\overline{I_{b}}$ and $A(a,b)=0$ otherwise. A finite (or infinite) word $\mathbf{a}=a_{1}\cdots a_{n}$ is admissible if $A(a_{i},a_{i+1})=1$ for all $i$. For each admissible pair $ab$ we write $g_{ab}:I_{b}\to I_{a}$ for the inverse branch; the collection $\Phi=\{g_{ab}:ab\in\mathcal{A}^{2}\ \text{admissible}\}$ is a (Markov) IFS. The symbolic space is $\Sigma^{\mathbb{N}}:=\{\mathbf{a}\in\mathcal{A}^{\mathbb{N}}\mid\forall j\in\mathbb{N}:A(a_{j},a_{j+1})=1\}$; we denote $\Sigma^{n}$ the set of admissible words of length $n$ and $\Sigma^{*}=\bigcup_{n\geq 1}\Sigma^{n}$. We define the notation $\mathbf{a}\to x$ to simply mean that $g_{\mathbf{a}}$ is defined at $x$. By denoting the last letter of $\mathbf{a}$ with $b(\mathbf{a})$, the condition $\mathbf{a}\to x$ satisfies $\mathbf{a}\to x\Longleftrightarrow x\in I_{b\left(\mathbf{a}\right)}$.
We assume:
  • (1)
    Uniform $\kappa_{+}$-contraction: there exists $\kappa_{+}\in(0,1)$ such that for every admissible $ab\in\Sigma^{2}$ there is $\kappa_{-}(ab)\in(0,1)$ with
    $\kappa_{-}(ab)\leq|g_{ab}^{\prime}(x)|\leq\kappa_{+}\quad\text{for all }x\in I_{b}.$
  • (2)
    Bounded $(C,\alpha)$-distortion: there exist $C>0$ and $\alpha>0$ such that, for all admissible $ab$ and all $x\neq y\in I_{b}$,
    $\big|\log|g_{ab}^{\prime}(x)|-\log|g_{ab}^{\prime}(y)|\big|\leq C|x-y|^{\alpha}.$
We say that $\Phi$ is $C^{1+\alpha}$ regular. When $\Phi$ is $C^{2}$, the bounded $(C,1)$-distortion follows from the Adler condition $\big\|\tfrac{T^{\prime\prime}}{(T^{\prime})^{2}}\big\|_{\infty}<\infty$ by the mean value theorem. Uniform contraction guarantees the existence of a unique attractor $F\subset I$, namely
$F=F_{\Phi}=\bigcap_{n\geq 1}\ \bigcup_{\mathbf{a}\in\Sigma^{n}}\overline{I_{\mathbf{a}}},\qquad I_{\mathbf{a}}:=g_{\mathbf{a}}(I_{b(\mathbf{a})}).$
Gibbs measures and equilibrium states. We work with Gibbs measures $\mu_{\phi}$ supported on $F$. A function $\phi:F\to\mathbb{R}$ has summable variations if
$\sum_{n\geq 2}\ \sup_{\mathbf{a}\in\Sigma^{n}}\ \sup_{x,y\in I_{\mathbf{a}}}|\phi(x)-\phi(y)|<\infty.$
A potential $\phi$ with summable variations is normalized if $\phi<0$ and $1$ is an eigenfunction of the transfer operator
$\mathcal{L}_{\phi}f(x):=\sum_{a\in\mathcal{A},\,ab\in\Sigma^{2}}e^{\phi(g_{ab}(x))}\,f(g_{ab}(x)),\qquad x\in I_{b},$
for all continuous $f:F\to\mathbb{C}$; we write $w_{ab}(x):=e^{\phi(g_{ab}(x))}$. The associated Gibbs measure $\mu_{\phi}$ is the unique probability measure on $F$ satisfying the $\phi$-conformality:
for all continuous $f$. In this setting $\mu_{\phi}$ is non-atomic and $T:F\to F$ is ergodic with respect to $\mu_{\phi}$.
Light tail (infinite alphabet). To cover parabolic systems (hence infinite alphabets), we prevent $\mu_{\phi}$ from concentrating on words with exceptionally small derivatives:
  • (3)
    Light $(C_{\phi,\gamma_{0}},\gamma_{0})$-tail: for some $\gamma_{0}>0$,
    $C_{\phi,\gamma_{0}}:=\sum_{ab\in\Sigma^{2}}\ \sup_{x\in I_{b}}e^{\phi(g_{ab}(x))}\,|g_{ab}^{\prime}(x)|^{-\gamma_{0}}<\infty.$
Non-linearity and temporal distance. The key input for Fourier decay (when the $g_{\mathbf{a}}^{\prime}$ are not constant) is a statistical non-concentration property for the distribution of derivatives along the symbolic tree. For $\mathbf{a},\mathbf{b}\in\Sigma^{n}$ and $x,y\in F$, define
$\Delta_{n}\big((\mathbf{a},x),(\mathbf{b},y)\big):=\big(\log g_{\mathbf{a}}^{\prime}(x)-\log g_{\mathbf{b}}^{\prime}(x)\big)-\big(\log g_{\mathbf{a}}^{\prime}(y)-\log g_{\mathbf{b}}^{\prime}(y)\big),$
which is a finite-time analogue of the temporal distance function associated to the roof $\tau:=-\log T^{\prime}$:
$\Delta_{\infty}\big((w,x),(v,y)\big):=\sum_{n=0}^{\infty}\Big(\tau(g_{w_{1}\cdots w_{n}}x)-\tau(g_{v_{1}\cdots v_{n}}x)\Big)-\Big(\tau(g_{w_{1}\cdots w_{n}}y)-\tau(g_{v_{1}\cdots v_{n}}y)\Big),$
so that $\Delta_{\infty}((w,x),(v,y))=\lim_{n\to\infty}\Delta_{n}((w|n,x),(v|n,y))$. In the flow picture (e.g. [44]), fluctuations of temporal distance govern mixing rates; in the present IFS setting the words $\mathbf{a},\mathbf{b}$ mimic the stable direction while $x,y$ play the role of the unstable direction under the expansion of $T$. To capture non-linearity at low regularity we couple motion in both symbolic and geometric directions. We assume the following multiscale incidence bound we call Quantitative Non-Linearity (QNL): there exist $\rho,\Theta\in(0,1)$ such that for every interval $I\subset\mathbb{R}$ and $n\in\mathbb{N}$,
$\sum_{\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}\in\Sigma^{n}}\mu(I_{\mathbf{a}})\mu(I_{\mathbf{b}})\mu(I_{\mathbf{c}})\mu(I_{\mathbf{d}})\,\mathbf{1}_{I}\!\left(\Delta_{n}\big((\mathbf{a},x_{\mathbf{c}}),(\mathbf{b},x_{\mathbf{d}})\big)\right)\ \lesssim\ |I|^{\Theta}+\rho^{n},$
where $x_{\mathbf{a}},x_{\mathbf{b}},x_{\mathbf{c}},x_{\mathbf{d}}$ denote the centers of $I_{\mathbf{a}},I_{\mathbf{b}},I_{\mathbf{c}},I_{\mathbf{d}}$. Under QNL we obtain power decay for oscillatory integrals with slowly varying amplitudes and phases.
@@ 1.2. Results in $C^{1+\alpha}$ regularity
Let $\Phi$ be a $C^{1+\alpha}$ IFS satisfying (1) uniform $\kappa_{+}$-contraction, (2) bounded $(C,\alpha)$-distortion, and let $\phi$ be a normalized potential satisfying the light-tail condition (3). Let $\mu_{\phi}$ be the associated Gibbs measure. If (QNL) holds, then there exists $\beta>0$ such that
@@
Theorem 1.1 follows ultimately similar lines as the works of Bourgain-Dyatlov [13] and several subsequent works [30, 41, 27, 10, 9] in higher generality. However, the multiscale QNL condition in Theorem 1.1 is in general very difficult to verify, it requires often quite delicate non-concentration analysis, Diophantine or spectral gap assumptions (see e.g. [40] for details and historical aspects), and these themselves are the main culprit that require smoothness from the dynamics. The main novelty of our work, which we discuss now, comes from the induction-on-scales scheme that allows us to prove QNL condition for systems that also involve truly $C^{1+\alpha}$ systems and in smoother categories completely avoiding the use of spectral gaps of twisted transfer operators.
First of all, in the $C^{2}$ category (finite or infinite branch), prior works such as [41, 27, 10, 5] employ the following (UNI) condition going back to Chernov and Dolgopyat [21]:
  • (UNI)
    $|X_{\mathbf{a}}(x,y)-X_{\mathbf{b}}(x,y)|\ \geq\ c_{0}\,|x-y|^{\alpha},$
where
$X_{\mathbf{a}}(x,y):=S_{n}\tau\circ g_{\mathbf{a}}(x)-S_{n}\tau\circ g_{\mathbf{a}}(y)=\log g_{\mathbf{a}}^{\prime}(x)-\log g_{\mathbf{a}}^{\prime}(y),\qquad\mathbf{a}\to x,y,$
measures the spatial distortions of $g_{\mathbf{a}}$. (Typically checked with $\alpha=1$, but the arguments extend formally to $\alpha\in(0,1]$.) Under (UNI) we give a proof of QNL:
@@
If the IFS has a finite or countable set of branches, then (UNI) implies (QNL).
In the finite-branch case, Theorem 1.2 appears, e.g., in [41, 10, 30, 27, 5], via spectral gaps for twisted transfer operators derived from Dolgopyat estimates. Here to showcase the method also in smoother category, we give a proof avoiding Dolgopyat’s estimate, instead building a $C^{\alpha}$ version of the tree strategy previously employed by the first author in the smoother setting [28] using an induction-on-scales scheme. Theorem 1.2 is new for parabolic IFSs and underpins our later applications to Patterson–Sullivan measures in the parabolic case and intermittent systems such as Manneville–Pommeau maps, where UNI holds. We also note that for analytic (finite-branch) IFSs, power Fourier decay is already guaranteed by the existence of a branch with non-vanishing second derivative, by combining [5, 10, 1, 8] (though only for the specific phase $\psi(x,\xi)=x\xi$).
While $(c_{0},\alpha)$–(UNI) is formally stated for $C^{1+\alpha}$, verifying a uniform lower bound of order $|x-y|^{\alpha}$ on the entire attractor is hard without at least $C^{2}$ smoothness. For truly $C^{1+\alpha}$ systems one needs a condition that leverages the fractal geometry of $F_{\Phi}$ and couples symbolic and geometric directions (words $\mathbf{a},\mathbf{b}$ and points $x,y$). This is inspired by the work of Tsujii and Zhang [44], where the unstable and stable directions are coupled to capture the non-linearity. However, the main difference with [44, 26] is that in our setting there is no smooth dynamics we can rely on. We therefore introduce:
  • (MNL)
    Measure Non-Linearity for $\Phi$: Let $\delta:=\dim_{H}F_{\Phi}$, let $\mu$ be the equilibrium state for the potential $\delta\tau$, and assume $\tau\in C^{\delta}$. There exists $C_{\text{(MNL)}}$ such that, for all $N\geq 1$, there are $\mathbf{a},\mathbf{b}\in\mathcal{A}^{N}$ with
    $\forall x_{0}\in F_{\Phi},\ \forall I\subset\mathbb{R},\quad\big(X_{\mathbf{a}}(\cdot,x_{0})-X_{\mathbf{b}}(\cdot,x_{0})\big)_{*}\mu(I)\ \leq\ C_{\text{(MNL)}}\,|I|.$
Using an induction-on-scales scheme, we then show that MNL indeed yields QNL in the finite-branch setting:

### v3

# arXiv:2505.15468v3[math.DS] 01 Mar 2026

@@ Abstract.
We establish power Fourier decay for equilibrium states of parabolic $C^{1+\alpha}$ iterated function systems with overlaps satisfying a multiscale nonlinearity condition. This class includes the Lyons conductance measures $\nu_{t}$, $0<t<1$, associated to Galton-Watson trees with equal weights yielding advance towards a conjecture of Lyons on the absolute continuity of $\nu_{t}$ for small $t$. Further applications include Patterson-Sullivan measures for cusped hyperbolic surfaces, extending the work of Bourgain and Dyatlov to parabolic settings, conformal measures for Manneville-Pommeau and Lorenz-type maps, and the construction of the first genuinely $C^{1+\alpha}$ IFSs whose attractors have positive Fourier dimension but are not $C^{1}$-conjugate to linear IFSs. The proof combines the Bourgain-Dyatlov sum-product strategy with a multiscale induction approach that bypasses the use of spectral gaps for twisted transfer operators needed in several other works in the area.
@@ 1. Introduction
1.1. Summary of the main results
There has been substantial recent interest in the high-frequency behaviour of Fourier transforms of measures $\widehat{\mu}(\xi)=\int e^{-2\pi i\xi x}\,d\mu(x)$, arising from random constructions [32, 71, 35, 21, 22, 41, 45], dynamical systems and quantum chaos [39, 38, 18, 47, 58, 2, 6, 67, 10, 13, 43, 44, 11, 42, 9], and additive combinatorics and geometric measure theory [59, 61, 28, 78, 40, 34, 3, 4]; see also [66]. The study of power Fourier decay is a signature of smoothness of the measure, and can be used to study equidistribution and Diophantine properties of numbers in the support [26, 64] and it is also closely related to proving absolute continuity notably in Shmerkin’s work on Bernoulli convolutions [70] and higher dimensions by Lindenstrauss and Varjú [49].
For dynamically defined measures, most literature concern uniformly hyperbolic systems with $C^{2}$ or analytic regularity and separation, where Fourier decay follows from spectral gaps for twisted transfer operators via Dolgopyat’s method [38, 18, 58, 2, 6, 67, 10, 13, 43, 44]. In contrast, many natural measures arise from overlapping systems (Bernoulli convolutions [70], Axiom $A$ diffeomorphisms [42]) or from parabolic dynamics (for example Lorenz-type systems [52]), and fall outside this framework. An illustrative example is the family of Lyons conductance measures $\nu_{t}$, $0<t<1$, describing effective conductance in Galton-Watson trees with equal weights [53] arising as self-conformal measures for nonlinear and overlapping parabolic IFSs. Their smoothness properties such as absolute continuity remain difficult to access [72, 57] and as with Bernoulli convolutions they are in fact conjectured to be absolutely continuous for all small $t$ by Lyons [53].
The purpose of this paper is to introduce a purely symbolic multiscale induction strategy to prove power Fourier decay for equilibrium states of overlapping parabolic $C^{1+\alpha}$ iterated function systems under a Quantitative Non-Linearity condition (Theorem 1.1), for which we provide broadly applicable criteria (Theorem 1.2). Our approach completely bypasses spectral gaps for twisted transfer operators [46, 67, 7, 13, 11], allowing us to treat genuinely $C^{1+\alpha}$ and parabolic systems beyond the scope of current smooth hyperbolic methods. As an application, we prove power Fourier decay for the Lyons conductance measures $\nu_{t}$ for all $0<t<1$ (Theorem 1.3). We note the analogous problem for Bernoulli convolutions would be useful for proving absolute continuity [70] and our results highlights the strong effect of nonlinearity as an amplifying effect for smoothness.
Further applications of Theorem 1.1 include power Fourier decay of Patterson–Sullivan measures for hyperbolic surfaces with parabolic elements, extending Bourgain–Dyatlov who used these in the study of scattering resonances in the convex co-compact setting [18] (Theorem 1.4), and all $\delta$-conformal measures for Manneville–Pomeau [65] and Lorenz-type maps arising from the Lorenz system [52] (Theorem 1.5). The multiscale induction we develop is flexible enough to treat genuinely $C^{1+\alpha}$ systems where smooth tools are unavailable. In particular, Theorem 1.1 yields the first examples of $C^{1+\alpha}$ IFSs not conjugate to linear systems whose equilibrium measures exhibit power Fourier decay under a condition we call Measure Non-Linearity (Theorem 1.6).
We now proceed to state the main Fourier decay theorem under QNL in the next section and criteria to check QNL, and then discuss in detail the applications above.
1.2. QNL and power Fourier decay
Let $\mathcal{A}$ be a finite or countable alphabet and let $\Sigma^{\mathbb{N}}=\{(a_{1},a_{2},\ldots)\in\mathcal{A}^{\mathbb{N}}:\ A(a_{i},a_{i+1})=1\}$ be a subshift of finite type with admissible words $\Sigma^{n}$ and $\Sigma^{*}=\bigcup_{n\geq 1}\Sigma^{n}$. Let $\{I_{a}\}_{a\in\mathcal{A}}$ be open intervals in $[0,1]$ and for each admissible pair $ab$ let $g_{ab}:I_{b}\to I_{a}$ be a $C^{1+\alpha}$-diffeomorphic contraction, defining $g_{\mathbf{a}}:=g_{a_{1}a_{2}}\circ\cdots\circ g_{a_{n-1}a_{n}}$ for $\mathbf{a}\in\Sigma^{n}$. Assume uniform $\kappa_{+}$-contraction $\kappa_{-}(ab)\leq|g_{ab}^{\prime}(x)|\leq\kappa_{+}<1$ and bounded $(C,\alpha)$-distortion $|\log|g_{ab}^{\prime}(x)|-\log|g_{ab}^{\prime}(y)||\leq C|x-y|^{\alpha}$ for all $x,y$. Then there exists attractor $F=F_{\Phi}:=\bigcap_{n\geq 1}\ \bigcup_{\mathbf{a}\in\Sigma^{n}}\overline{g_{\mathbf{a}}(I_{a_{n}})}$. Let $\phi$ be a normalised potential with summable variations and let $\mu_{\phi}$ be the associated Gibbs measure. In the countable-branch case we additionally assume the light-tail condition: there exists $\gamma_{0}>0$ such that
$\sum_{ab\in\Sigma^{2}}\ \sup_{x\in I_{b}}e^{\phi(g_{ab}(x))}\,|g_{ab}^{\prime}(x)|^{-\gamma_{0}}<\infty.$
For $\mathbf{a},\mathbf{b}\in\Sigma^{n}$ and $x,y\in F$ define
$\Delta_{n}((\mathbf{a},x),(\mathbf{b},y)):=(\log g_{\mathbf{a}}^{\prime}(x)-\log g_{\mathbf{b}}^{\prime}(x))-(\log g_{\mathbf{a}}^{\prime}(y)-\log g_{\mathbf{b}}^{\prime}(y)).$
We say that $\Phi$ satisfies (QNL) if there exist $\rho,\Theta\in(0,1)$ such that for any interval $I\subset\mathbb{R}$ and $n\in\mathbb{N}$,
$\sum_{\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}\in\Sigma^{n}}\mu(I_{\mathbf{a}})\mu(I_{\mathbf{b}})\mu(I_{\mathbf{c}})\mu(I_{\mathbf{d}})\,\mathbf{1}_{I}\!\left(\Delta_{n}((\mathbf{a},x_{\mathbf{c}}),(\mathbf{b},x_{\mathbf{d}}))\right)\lesssim|I|^{\Theta}+\rho^{n}.$
@@ 1.2. QNL and power Fourier decay
Let $\Phi$ be a $C^{1+\alpha}$ IFS as above, satisfying uniform $\kappa_{+}$-contraction and bounded $(C,\alpha)$-distortion. Let $\phi$ be a normalised potential with summable variations satisfying the light-tail condition, and let $\mu_{\phi}$ be the associated Gibbs measure without atoms. If (QNL) holds, then there exists $\beta>0$ such that
@@
To make Theorem 1.1 usable in applications, we seek checkable sufficient conditions for (QNL). In the $C^{2}$ uniformly hyperbolic setting, Fourier decay is typically obtained by establishing spectral gaps for twisted transfer operators via Dolgopyat-type estimates (e.g. [47, 67, 7, 13, 11]). A dynamical formulation of the required non-concentration is the Uniform Non-Integrability (UNI) condition introduced by Chernov and Dolgopyat [31]. Defining cocycle $\tau(g_{\mathbf{a}}(x)):=-\log|g_{\mathbf{a}}^{\prime}(x)|$ and fluctuation $X_{\mathbf{a}}(x,y):=\tau(g_{\mathbf{a}}(x))-\tau(g_{\mathbf{a}}(y))$, we say that $\Phi$ satisfies (UNI) if there exist $c_{0}>0$ and infinite $N$ such that for every $c\in\mathcal{A}$ there exist $\mathbf{a},\mathbf{b}\in\Sigma^{N}$ with $b(\mathbf{a})=b(\mathbf{b})=c$ satisfying
$|X_{\mathbf{a}}(x,y)-X_{\mathbf{b}}(x,y)|\geq c_{0}|x-y|^{\alpha}\quad\text{for all }x\neq y\in F\cap I_{c}.$
However, in genuinely $C^{1+\alpha}$ setting verifying UNI on the entire attractor becomes difficult leading to many difficulties e.g. in smooth dynamics [76, 42]. In these cases one instead seeks a statistical form of non-concentration along the equilibrium state. Letting $\delta=\dim_{H}F$ be the Hausdorff dimension and $\mu$ be the equilibrium state for $\delta\tau$, we say that the IFS $\Phi$ satisfies the Measure Non-Linearity (MNL) condition if there exists $C_{\mathrm{(MNL)}}>0$ such that for all $N$ there are $\mathbf{a},\mathbf{b}\in\Sigma^{N}$ with
$\forall x_{0}\in F,\ \forall I\subset\mathbb{R},\qquad\big(X_{\mathbf{a}}(\cdot,x_{0})-X_{\mathbf{b}}(\cdot,x_{0})\big)_{*}\mu(I)\ \leq\ C_{\mathrm{(MNL)}}\,|I|.$
This replaces the uniform lower bound in (UNI) by a measure-theoretic non-concentration property along the symbolic tree, in the spirit of the coupling of geometric and symbolic directions in [76].
Under UNI or MNL, we then have the needed QNL for Theorem 1.1:
@@
If infinite $C^{1+\alpha}$ IFS has (UNI), or in finite branch case (MNL), then (QNL) holds.
Note here we allow arbitrary overlaps. The UNI part in finite branch case can be verified using spectral gaps for twisted transfer operators as in [67] and in overlapping case a co-cycle version of the Dolgopyat method was needed and indeed developed in [7, 13] via average contraction of random Dolgopyat operators. Our strategy avoids these constructions and replaces it by the multiscale induction allowing us to access parabolic and $C^{1+\alpha}$ cases.
Let us now discuss applications of Theorem 1.1 and Theorem 1.2 reducing power Fourier decay to UNI or MNL type statements.
1.3. Smoothness of effective conductance of Galton-Watson trees.
Let $(p_{k})_{k\geq 0}$ be a probability distribution on $\mathbb{N}\cup\{0\}$ and let $T$ be the random genealogical tree obtained from the associated Galton-Watson branching process begun with one individual $\varnothing$, in which each vertex has $k$ children with probability $p_{k}$, independently of all other vertices. Adjoining a parent vertex $v$ to $\varnothing$ and connecting it to $\varnothing$ by an edge, we may view the resulting graph $T\cup\{v\}$ as an electrical network in which each edge has unit conductance. Writing $\Theta$ for the effective conductance of $T\cup\{\mathbf{v}\}$ from $v$ to infinity, we know that $\Theta$ equals to the probability of a simple random walk starting from $v$ and returning to $v$ [55, 54, 25]. Then the effective conductance is a random continued fraction $\Theta=\frac{1}{1+\frac{1}{X_{1}+\frac{1}{1+\cdots}}}$ for i.i.d. $X_{1},X_{2},\ldots$. In [53], Lyons considered the special case where $X_{i}\in\{0,t\}$ i.i.d. with equal probabilities. Thus in this case $\Theta$ is obtained as the infinite random composition $\Theta=\lim_{n\to\infty}f_{X_{1}}\circ f_{X_{2}}\circ\cdots f_{X_{n}}(0)$ for $f_{0}(x)=\frac{x}{1+x}$ and $f_{t}(x)=\frac{x+t}{1+x+t}$ so if we denote by $\nu_{t}$ the law of $\Theta$, it is an equilibrium state for the overlapping IFS $\{f_{0},f_{t}\}$, which is parabolic as $f_{0}^{\prime}(0)=1$.
On the smoothness of $\nu_{t}$, Lyons [53] proved $\nu_{t}$ is singular for any $t>t_{0}\approx 0.2689$ and conjectured $\nu_{t}$ should to be absolutely continuous for all small $t$. This was motivated by numerics and it being a nonlinear analogue of the Bernoulli convolution $\mu_{\lambda}$ (distribution of $\sum\pm\lambda^{n}$), where a similar conjecture is widely open despite much recent progress [70, 37, 77] and for higher dimensional analogues of Bernoulli convolutions [50]. As with Bernoulli convolutions, absolute continuity and density information of $\nu_{t}$ are known for almost every $t$ by transversality method [72, 73, 75, 16]. In [70] Shmerkin related the absolute continuity of $\mu_{\lambda}$ to the power Fourier decay of Bernoulli convolutions associated to powers $\lambda^{k}$, but it remains a very challenging problem prove this, where the Erdös-Kahane method can be applied for almost every $\lambda$ (see e.g. [74, 70]).
By inducing the parabolic Lyons IFS [57], as a consequence of Theorem 1.1 we have: